# Hitting probability for Reflected Brownian Motion at Small Target

### v2

@@ Abstract.
We derive the asymptotic behavior of hitting probability at small target of size $O(\epsilon)$ for reflected Brownian motion in domains with suitable smooth boundary conditions, where the boundary of domain contains both reflecting part, absorbing part and target. In this case the domain could be localized near the target and explicit computations are possible. The asymptotic behavior is only related to $\epsilon$ up to some multiplicative constants that depends on the domain and boundary conditions.
@@ 1. Introduction
Reflected Brownian motions, or more generally reflected Wiener processes, can roughly be thought as a Brownian motion that is “reflected” when it hits some targets. Although it has simple intuitive description, it is extremely hard to construct Reflected Brownian motions in domains in $\mathbb{R}^{n}$ for $n\geq 2$ [1, 2]. Despite of this difficulty, reflected Brownian motion has became a great area of research in both pure mathematics and applied mathematics [1, 3, 4]. In particular, its hitting behavior at small targets has grasped the interests of more and more researchers [5, 6]. Despite there has been some works done in asymptotic hitting behaviors of Brownian motions at small targets [7, 8], the asymptotic hitting behaviors of reflected Brownian Motions at small targets is poorly understood.
@@
In this paper we will mainly be focused on asymptotic hitting probability for reflected Brownian motion at small targets. The current state of art in this area is preforming explicit computations using connections between reflected Brownian motion and partial differentiate equations, which was done by Denis S. Grebenkov and Adrien Chaigneau in certain domains [4]. However, this approach is limited to huge amount of computations and could only be done in a few domains where solutions to Dirichlet problem behaves nicely. Instead, in this paper we observed that the reflected Brownian motion behaves very well in a large class of domains which we defined to be “Smooth Uniform Lipchitz Domains” (see Definition 2.5), and in this class of domains the asymptotic behavior hitting probability is comparable to a fundamental solution of Laplacian. The intuition is when the reflect Brownian particle is near the target, the probability that it hits non-targeted boundary is very small and thus irrelevant to the universal shape of the domain. When target is very small, the local geometry of SULD domain near the target is nearly half space, so the asymptotic hitting probability should agree to the case of a half ball with target at the bottom planar surface, and computation could be done easily in this case.
@@
Complex analysis, especially conformal maps and harmonic measures, has played a crucial role in studies of planar Brownian motions [9, 10]. We will also use methods from complex analysis to prove a more accurate result (see Theorem 3.4). Regrettably, this accurate result haven’t be generalized to higher dimensions as the method of complex analysis ceases to work.
@@
Although this paper successfully computed the asymptotic behavior of hitting probability in a large class of domains, the asymptotic mean hitting time still remains open except for some elongated domains [6]. Besides, the asymptotic hitting probability for domains with rough boundaries (e.g. fractal domains) still remains open. The methods in this paper all cease to work for these kind of domains and even some basic properties of reflected Brownian motion hasn’t been well-studied for these domains yet.
@@
In this paper, we will construct reflected Brownian motion in chapter 2, and show any analytic simply connected domain in $\mathbb{C}$ is SULD. Some technique proofs will be included in chapter 7 appendix. In chapter 3 we define the target, absorbing boundaries and state our main results (Theorem 3.3 and Theorem 3.4). In chapter 4 we cite some standard definitions and results in complex analysis. We will prove Theorem 3.3 and Theorem 3.4 using complex analysis tools in $\mathbb{C}$ in chapter 5 and prove Theorem 3.3 in higher dimension in chapter 6.
@@ 2. Domains and Construction of Reflected Brownian Motion
In this section we construct the reflected Brownian motions (RBM) in certain domains.
@@ 2.1. Construction of RBM on General (SULD) Domains
In this subsection we discuss how to define RBM on any connected domains in $\mathbb{R}^{n}$ $(n\geq 2)$ with suitable smooth and analytical conditions.
@@
There are many equivalent ways to construct RBM, in here we follow the construction by [2]. Constructions in more general domain could be done, for example, in [11] using Dirichlet form. We define the RBM by explicit construction of solution to stochastic differential equation which generate RBM in an intuitive sense. Let $\mathbb{H}_{n}:=\mathbb{R}^{n}_{+}$ be the upper half space in $\mathbb{R}^{n}$ and $W_{t}$ be a $n$-dimensional unrestricted standard Brownian motion. Consider the stochastic differential equation in $(X_{t}^{x},\xi_{t}^{x})$:
(1) $dX_{t}^{x}=\mathrm{d}W_{t}+\mathbf{1}_{\partial\mathbb{H}_{n}}(X_{t}^{x})\gamma(X_{t}^{x})\mathrm{d}\xi_{t}^{x}$
with $X_{0}^{x}=x$ and $\xi_{0}^{x}=0$, and $\gamma$ is a constant vector valued function with $\gamma\equiv(1,0,...,0)$. $(X_{t}^{x},\xi_{t}^{x})$ is adapted to the $\sigma$-algebra that with probability $1$, $\xi_{t}^{x}$ is non-decreasing in $t$ and increases only at set $\Delta:=\{t:X^{x}_{t}\in\partial\mathbb{H}_{n}\}$ and $\Delta$ is a Lebesgue null set. Now we use explicit construction to proof the following proposition
@@
Under the condition above, the stochastic differential equation (1) has a unique solution pair $(X_{t}^{x},\xi_{t}^{x})$ in the sense that any other pair satisfies (1) and $\Delta$ be a Lebesgue null set is equal to this pair with probability 1. We define the reflected Brownian motion in +$\mathbb{R}^{n}_{+}$ to be the process $X^{x}_{t}$.[2]
@@
Let $X^{x}_{t}$ be defined in 2.2. Let $Y^{x}_{t}=(Y^{x}_{1}(t),Y^{x}_{2}(t),...,Y^{x}_{n}(t))$ be the standard Brownian motion in $R^{n}$ such that $Y^{x}_{0}=x=(x_{1},x_{2},...,x_{n})\in\mathbb{H}_{n}$ and each $Y_{k}^{x}(t)$ is the independent one dimensional Brownian motion. Define the process
@@
Then $(Y^{x}_{t},L^{x}_{t})$ is equal to $(X_{t}^{x},\xi_{t}^{x})$ constructed in Proposition 2.2 with probability 1.
Proof.
By Tanaka’s formula we have
[…]
where $Z^{x}_{1}$ is a Brownian motion. Thus $(Y^{x}_{t},L^{x}_{t})$ satisfies (1). As Brownian motion spend almost zero time on any point set we see that $\Delta=\{t:Y^{x}_{1}(t)=0\}$ is a Lebesgue null set. So by Proposition 2.2 we see that $(Y^{x}_{t},L^{x}_{t})$ is equal to $(X_{t}^{x},\xi_{t}^{x})$ with probability 1. ∎
@@
(2) $dX_{t}^{x}=\mathrm{d}W_{t}+\mathbf{1}_{\partial\Omega}(X_{t}^{x})\gamma(X_{t}^{x})\mathrm{d}\xi_{t}^{x}$
has a unique solution pair $(X_{t}^{x},\xi_{t}^{x})$ in the sense that any other pair is equal to this pair with probability 1. In particular when $\gamma$ is the unit normal vector at each $x$, we define the reflected Brownian motion in $\Omega$ to be the corresponding solution process process $X^{x}_{t}$.[2]
@@
    If $k(x)=0$ only then the process is just the normal Brownian motion as boundary is not involved;
@@
    If we can choose $k(x)>0$ then the coordinate mapping $\Phi_{k}$ associated with $U^{k(x)}$ changes the problem to a problem in the upper half space with normal reflection on boundary. The new process $\Phi_{k}(X_{t})$ satisfies a new stochastic differential equation by itô’s formula: Write $\Phi_{k}=f=(f^{1},...,f^{n})$, then
    $\mathrm{d}f^{i}(X_{t})=\left(\sum^{n}_{k=1}\frac{\partial f^{i}}{\partial x_{k}}\mathrm{d}W^{k}_{t}\right)+\frac{1}{2}\Delta f^{i}\mathrm{d}t+\mathbf{1}_{\partial\Phi_{k}(U_{k})\cap\overline{\mathbb{H}}}(f(X_{t}))\delta_{i1}|\nabla f^{1}|\mathrm{d}\xi_{t}$
@@
    This problem was solved above and we obtain a process in the upper half space with normal reflection. Then $\Phi^{-1}_{k}$ maps the process in upper half space back to $U^{k(x)}$ which gives us $X_{t}^{x},0\leq t\leq S_{1}$. For this process $(1)$ will hold, with $\xi_{t}^{x}$ being identical to the local time on the boundary for the process in upper half space. Then we just repeat this and gives us the desired process.
However, non-uniqueness choices of $k(x)$ result in the same process. We will have to additionally show the consistency of our definition which is missed in [2].
@@
We give another equivalent construction of RBM in analytic simply connected domains in $\mathbb{C}$ as a special case in order to apply tools in complex analysis to proof a stronger result.
2.2. Equivalent Construction of RBM on Analytic Simply Connected Domains in $\mathbb{C}$
@@
Let $\Omega\subsetneq\mathbb{C}$ be a domain in the sense that it is open and connected. Let $\partial\Omega:=\bar{\Omega}\setminus\mathring{\Omega}$. We say $\Omega$ is analytic if there is an univalent function $f$ analytic near $\partial\mathbb{D}$ such that $\partial\Omega$ is the image of $\partial\mathbb{D}$ under $f$. We say $\Omega$ is analytic simply connected domain if it is additionally simply connected.
One equivalent construction of the RBM on analytic domain is using conformal mapping. We define RBM in upper half plane $\mathbb{H}$ then define the RBM in general analytic simply connected domain $\Omega$ by conformal maps, where Riemann mapping theorem ensure there is an conformal map between $\mathbb{H}$ and $\Omega$.
@@
Let $(X_{t},Y_{t})$ be the standard Brownian motion in $\mathbb{R}^{2}$ where $X_{t},Y_{t}$ are independent standard Brownian motion in $\mathbb{R}$. Define another stochastic process $B_{t}=(X_{t},|Y_{t}|)$ as the (complex) reflected Brownian motion in $\mathbb{H}$. Identify $\mathbb{C}$ with $\mathbb{R}^{2}$ our reflected Brownian motion in $\mathbb{H}$ is $B_{t}=X_{t}+i|Y_{t}|$. We write $B^{z}_{t}$ to denote the reflected Brownian motion starts at $z$. We also note the process $X_{t}+iY_{t}$ is referred to the complex Brownian motion by [9].
By Proposition 2.3 we see the stochastic process $B_{t}$ is equal to RBM in $\mathbb{H}$ constructed in proof of Proposition 2.2 with probability 1. We now construct RBM in unit disk $\mathbb{D}:=\{z\in\mathbb{C}:|z|<1\}$.
@@
By the above proof and Proposition 2.6 we can construct the RBM in $\mathbb{D}$. Alternatively we have another equivalent construction of RBM in $\mathbb{D}$. We firstly prove a powerful characterisation of conformal invariance of RBM in $\mathbb{H}$.
@@
Suppose $\Omega$ and $\Omega^{\prime}$ are two simply connected open subset of $\mathbb{H}$ and $B^{z_{0}}_{t}$ is the RBM in $\mathbb{H}$ where $z_{0}\in\Omega$ with the associated local time process $\xi^{z_{0}}_{t}$. Let $f$ be a conformal automorphism of $\mathbb{H}$. Let $\tau_{\Omega}=\inf\{t\geq 0:B^{z_{0}}_{t}\in\partial\Omega\}$. Then with probability one, $(f(B^{z_{0}}_{t}),t\in[0,\tau))$ is equal to the time-changed RBM in $\mathbb{H}$.
@@
By examine above proof and notice that any conformal map $f=u+iv$ from $\Omega$ to $\mathbb{H}$ that can be extended conformally between a neighborhood of $\Omega$ and $\mathbb{H}$ with finite number of singularities (as we may delete them as the probability that the complex Brownian motion hit them is 0) on $\partial\Omega$, we have a stronger proposition:
@@
Let $B^{z}_{t}$ be defined in Definition 2.9. Let ${f(z)=i\frac{z-i}{z+i}}$. Then the process $f(B^{z}_{t})$ is the RBM started at $f(z)$ which is equivalent to complex RBM in $\mathbb{D}$ constructed by Proposition 2.6 after a suitable time change, which is denoted by $X^{f(z)}_{t}$.
Proof.
We notice that the point $\infty$ is mapped to $i$ by $f$, which could be removed from consideration as $\{i\}$ is a null set. We note that by above proposition $(f^{j})^{-1}\circ f$ is an automorphism of $\mathbb{H}$ and thus it maps $B^{z}_{t}$ to another time-changed complex RBM in $\mathbb{H}$. It follows that $f(z)$ is the time changed complex RBM in $\mathbb{D}$. ∎
By above arguments we can now give an equivalent definition of the complex RBM via conformal mapping:
@@
Suppose $B_{t}$ is the complex RBM in $\mathbb{D}$ and let $\Omega$ be an analytic simply connected domain. Let $g$ be the associated conformal map from $\Omega$ to $\mathbb{D}$ above. Then we define the process $\tilde{B_{t}}=g^{-1}(B_{t})$ to be the reflected Brownian motion in $\Omega$.
Proof.
By lemma 7.3 we see that $f$ could be extended so $f\circ(f^{i})^{-1}$ could be the associated maps (properties could be checked by procedure similarly to proof of proposition 2.10 and by open mapping theorem each open cover of $\mathbb{D}$ is mapped to open cover of $\Omega$ by $f$. ∎
We would like to finish this section by the following key proposition, which is a generalization of proposition 2.12. Its proof is also similar to the proof we did above.
@@
Let $\Omega_{1}$ and $\Omega_{2}$ be two analytic simply connected domain. By Riemann mapping theorem there is a conformal map $f$ between $\Omega_{1}$ and $\Omega_{2}$ which could also be extended to a conformal map between open neighborhood of $\overline{\Omega_{1}}$ and $\overline{\Omega_{2}}$. Let $B_{t}$ be the complex RBM in $\Omega_{1}$, then $f(B_{t})$ is the time-changed complex RBM in $\Omega_{2}$, starts at $f(z_{0})$.

### v4

@@ Abstract.
We derive the asymptotic behavior of hitting probability at a small target of size $O(\epsilon)$ for reflected Brownian motion in domains with suitable smooth boundary conditions, where the boundary of the domain contains both reflecting parts, absorbing part, and target. In this case, the domain could be localized near the target, making explicit computations possible. The asymptotic behavior is only related to $\epsilon$ up to some multiplicative constants that depend on the domain and boundary conditions.
@@ 1. Introduction
Reflected Brownian motions, or more generally reflected Wiener processes, can roughly be thought of as a Brownian motion that is “reflected” when it hits some subsets. Although they have simple, intuitive descriptions, it is extremely hard to construct Reflected Brownian motions in domains in $\mathbb{R}^{n}$ for $n\geq 2$ [1, 2]. Despite this difficulty, reflected Brownian motion has become a great area of research in both pure mathematics and applied mathematics [1, 3, 4]. In particular, its hitting behavior at small targets has grasped the interest of more and more researchers [5, 6]. Despite there has been some work done in asymptotic hitting behaviors of Brownian motions at small targets [7, 8], the asymptotic hitting behaviors of reflected Brownian Motions at small targets are poorly understood.
@@
In this paper, we will mainly be focused on asymptotic hitting probability for reflected Brownian motion at small targets. The current state of the art in this area is performing explicit computations using connections between reflected Brownian motion and partial differentiate equations, which was done by Denis S. Grebenkov and Adrien Chaigneau in certain domains [4]. However, this approach is limited to many computations and could only be done in a few domains where solutions to the Dirichlet problem behave nicely. Instead, in this paper, we observed that the reflected Brownian motion behaves very well in a large class of domains, which we defined to be “Smooth Uniform Lipchitz Domains” (see Definition Nicedomain), and in this class of domains, the asymptotic behavior hitting probability is comparable to a fundamental solution of Laplacian. The intuition is when the reflected Brownian particle is near the target, the probability that it hits the non-targeted boundary is very small and thus irrelevant to the universal shape of the domain. When the target is very small, the local geometry of the SULD domain near the target is nearly half-space, so the asymptotic hitting probability should agree to the case of a half ball with the target at the bottom planar surface, and computation could be done easily in this case.
@@
Complex analysis, especially conformal maps and harmonic measures, has played a crucial role in studies of planar Brownian motions [9, 10]. We will also use methods from complex analysis to prove a more accurate result (see Theorem 3.4). Regrettably, this accurate result hasn’t been generalized to higher dimensions as the complex analysis method ceases to work.
@@
Although this paper successfully computed the asymptotic behavior of hitting probability in a large class of domains, the asymptotic mean hitting time still remains open except for some elongated domains [6]. Besides, the asymptotic hitting probability for domains with rough boundaries (e.g. fractal domains) still remains open. The methods in this paper all cease to work for these kind of domains, and even some basic properties of reflected Brownian motion haven’t been well-studied for these domains yet.
@@
In this paper, we will construct reflected Brownian motion in Section 2, and show any analytic simply connected domain in $\mathbb{C}$ is SULD. Some technique proofs will be included in Section 7 appendix. In Section 3, we define the target, absorbing boundaries, and state our main results (Theorem 3.3 and Theorem 3.4). In Section 4, we cite some standard definitions and results in complex analysis. We will prove Theorem 3.3 and Theorem 3.4 using complex analysis tools in $\mathbb{C}$ in Section 5 and prove Theorem 3.3 in the higher dimensions in Section 6.
@@ 2. Domains and Construction of Reflected Brownian Motion
In this section, we construct the reflected Brownian motions (RBM) in certain domains.
@@ 2.1. Construction of RBM on General (SULD) Domains
In this subsection, we discuss how to define RBM on any connected domains in $\mathbb{R}^{n}$ $(n\geq 2)$ with suitable smooth and analytical conditions.
@@
There are many equivalent ways to construct RBM. Here we follow the construction by [2]. Constructions in more general domain could be done, for example, in [11] using Dirichlet form. We define RBM by explicitly constructing the solution to a stochastic differential equation, which generates RBM in an intuitive sense. Let $\mathbb{H}_{n}:=\mathbb{R}^{n}_{+}$ be the upper half space in $\mathbb{R}^{n}$ and $W_{t}$ be a $n$-dimensional unrestricted standard Brownian motion. Consider the stochastic differential equation in $(X_{t}^{x},\xi_{t}^{x})$:
(1) $\mathrm{d}X_{t}^{x}=\mathrm{d}W_{t}+\mathbf{1}_{\partial\mathbb{H}_{n}}(X_{t}^{x})\gamma(X_{t}^{x})\mathrm{d}\xi_{t}^{x}$
with $X_{0}^{x}=x$ and $\xi_{0}^{x}=0$, and $\gamma$ is a constant vector valued function with $\gamma\equiv(1,0,...,0)$. $(X_{t}^{x},\xi_{t}^{x})$ is adapted to the filtration that, with probability $1$, $\xi_{t}^{x}$ is non-decreasing in $t$ and increases only at set $\Delta:=\{t:X^{x}_{t}\in\partial\mathbb{H}_{n}\}$ and $\Delta$ is a Lebesgue null set. Now we use explicit construction to prove the following proposition
@@
Under the condition above, the stochastic differential equation (1) has a strong, pathwise unique solution pair $(X_{t}^{x},\xi_{t}^{x})$. We define the reflected Brownian motion in $\mathbb{H}_{n}$ to be the process $X^{x}_{t}$.[2]
@@
Let $X^{x}_{t}$ be defined in 2.2. Let $Y^{x}_{t}=(Y^{x}_{1}(t),Y^{x}_{2}(t),...,Y^{x}_{n}(t))$ be the standard Brownian motion in $R^{n}$ such that $Y^{x}_{0}=x=(x_{1},x_{2},...,x_{n})\in\mathbb{H}_{n}$ and each $Y_{k}^{x}(t)$ is the independent one-dimensional Brownian motion starts at $x$. Define the process
@@
$L_{t}^{x}=\frac{1}{2}\int_{0}^{t}\delta\left(Y^{x}_{1}(s)\right)\mathrm{d}s:=|Y^{x}_{1}(t)|-x-\int_{0}^{t}\operatorname{sign}\left(Y^{0}_{1}+x\right)\mathrm{d}Y^{0}_{1}$
where $Y^{x}_{1}=Y^{0}_{1}+x$ and $Y^{0}_{1}$ is a standard Brownian motion on $\mathbb{R}$. Then $(Y^{x}_{t},L^{x}_{t})$ is indistinguishable from $(X_{t}^{x},\xi_{t}^{x})$ constructed in Proposition 2.2.
Proof.
By Tanaka’s formula, we have
$|Y_{1}^{x}(t)|=|Y_{1}^{0}(t)+x|=x+\int_{0}^{t}\operatorname{sign}\left(Y^{x}_{1}\right)\mathrm{d}Y^{0}_{1}+L^{x}_{t}=Z^{x}_{1}(t)+L^{x}_{t}$
@@
(2) $\mathrm{d}X_{t}^{x}=\mathrm{d}W_{t}+\mathbf{1}_{\partial\Omega}(X_{t}^{x})\gamma(X_{t}^{x})\mathrm{d}\xi_{t}^{x}$
has a pathwise unique strong solution pair $(X_{t}^{x},\xi_{t}^{x})$. In particular when $\gamma$ is the unit normal vector at each $x$, we define the reflected Brownian motion in $\Omega$ to be the corresponding solution process process $X^{x}_{t}$.[2]
@@
    If $k(x)=0$ only then the process is just the normal Brownian motion as the boundary is not involved;
@@
    If we can choose $k(x)>0$, then the coordinate mapping $\Phi_{k}$ associated with $U^{k(x)}$ changes the problem to a problem in the upper half-space with normal reflection on the boundary. The new process $\Phi_{k}(X_{t})$ satisfies a new stochastic differential equation by itô’s formula: Write $\Phi_{k}=f=(f^{1},...,f^{n})$, then
    $\mathrm{d}f^{i}(X_{t})=\left(\sum^{n}_{k=1}\frac{\partial f^{i}}{\partial x_{k}}\mathrm{d}W^{k}_{t}\right)+\frac{1}{2}\Delta f^{i}\mathrm{d}t+\mathbf{1}_{\partial\Phi_{k}(U_{k})\cap\overline{\mathbb{H}}}(f(X_{t}))\delta_{i1}|\nabla f^{i}|\mathrm{d}\xi_{t}$
@@
    This problem was solved above, and we obtained a process in the upper half-space with normal reflection. Then $\Phi^{-1}_{k}$ maps the process in upper half space back to $U^{k(x)}$ which gives us $X_{t}^{x},0\leq t\leq S_{1}$. For this process $(1)$ will hold, with $\xi_{t}^{x}$ being identical to the local time on the boundary for the process in upper half space. Then we just repeat this, and it gives us the desired process.
However, non-uniqueness choices of $k(x)$ result in the same process. We will have to additionally show the consistency of our definition, which is missed in [2].
@@
We give another equivalent construction of RBM in analytic simply connected domains in $\mathbb{C}$ as a special case in order to apply tools in complex analysis to prove a stronger result.
2.2. Equivalent Construction of RBM in Domains in $\mathbb{C}$
@@
Let $\Omega\subsetneq\mathbb{C}$ be a domain in the sense that it is open and connected. Let $\partial\Omega:=\bar{\Omega}\setminus\mathring{\Omega}$. We say $\Omega$ is analytic if there is an univalent function $f$ analytic near $\partial\mathbb{D}$ such that $\partial\Omega$ is the image of $\partial\mathbb{D}$ under $f$. We say $\Omega$ is an analytic simply connected domain if it is additionally simply connected.
One equivalent construction of the RBM in the analytic domain is using conformal maps. We define RBM in the upper half plane $\mathbb{H}$ then define the RBM in general analytic simply connected domain $\Omega$ by conformal maps, where Riemann mapping theorem ensures there is a conformal map between $\mathbb{H}$ and $\Omega$.
@@
Let $(X_{t},Y_{t})$ be the standard Brownian motion in $\mathbb{R}^{2}$ where $X_{t},Y_{t}$ are independent standard Brownian motion in $\mathbb{R}$. Define another stochastic process $B_{t}=(X_{t},|Y_{t}|)$ as the (complex) reflected Brownian motion in $\mathbb{H}$. Identify $\mathbb{C}$ with $\mathbb{R}^{2}$ our reflected Brownian motion in $\mathbb{H}$ is $B_{t}=X_{t}+i|Y_{t}|$. We write $B^{z}_{t}$ to denote the reflected Brownian motion starting at $z$. We also note the process $X_{t}+iY_{t}$ is referred to the complex Brownian motion by [9].
By Proposition 2.3, we see the stochastic process $B_{t}$ is indistinguishable from the RBM in $\mathbb{H}$ constructed in proof of Proposition 2.2. We now construct RBM in unit disk $\mathbb{D}:=\{z\in\mathbb{C}:|z|<1\}$.
@@
By the above proof and Proposition 2.6, we can construct the RBM in $\mathbb{D}$. Alternatively, we have another equivalent construction of RBM in $\mathbb{D}$. We firstly prove a powerful characterization of conformal invariance of RBM in $\mathbb{H}$.
@@
Suppose $\Omega$ and $\Omega^{\prime}$ are two simply connected open subset of $\mathbb{H}$ and $B^{z_{0}}_{t}$ is the RBM in $\mathbb{H}$ where $z_{0}\in\Omega$ with the associated local time process $\xi^{z_{0}}_{t}$. Let $f$ be a conformal automorphism of $\mathbb{H}$. Then the process $f(B^{z_{0}}_{t})$ is indistinguishable from a time-changed RBM in $\mathbb{H}$.
@@
By examine above proof and notice that any conformal map $f=u+iv$ from $\Omega$ to $\mathbb{H}$ that can be extended conformally between a neighborhood of $\Omega$ and $\mathbb{H}$ with a finite number of singularities (as we may delete them as the probability that the complex Brownian motion hit them is 0) on $\partial\Omega$, we have a stronger proposition:
@@
Let $B^{z}_{t}$ be defined in Definition 2.9. Let ${f(z)=i\frac{z-i}{z+i}}$. Then the process $f(B^{z}_{t})$ is a time-changed RBM started at $f(z)$. In particular, $f(B^{z}_{t})$ is distinguishable from the time-changed complex RBM in $\mathbb{D}$ constructed by Proposition 2.6.
By the above arguments, we can now give an equivalent definition of the complex RBM via conformal mapping:
@@
Suppose $B_{t}$ is the complex RBM in $\mathbb{D}$ and let $\Omega$ be an analytic simply connected domain. Let $g$ be the associated conformal map from $\Omega$ to $\mathbb{D}$ above. Let
[…]
Then we define the process $\tilde{B_{t}}=g^{-1}(B_{\sigma(t)})$ to be the (complex) RBM in $\Omega$.
We would like to finish this section by the following key proposition, which is a generalization of a proposition 2.12. Its proof is also similar to the proof we did above.
@@
Let $\Omega_{1}$ and $\Omega_{2}$ be two analytic simply connected domains. By Riemann mapping theorem, there is a conformal map $f$ between $\Omega_{1}$ and $\Omega_{2}$ which could also be extended to a conformal map between the open neighborhoods of $\overline{\Omega_{1}}$ and $\overline{\Omega_{2}}$. Let $B_{t}$ be the complex RBM in $\Omega_{1}$, then $f(B_{t})$ is the time-changed complex RBM in $\Omega_{2}$, starts at $f(z_{0})$.